# ST-USleepNet: A Spatial-Temporal Coupling Prominence Network for Multi-Channel Sleep Staging

### v1

@@ Abstract
Sleep staging is critical for assessing sleep quality and diagnosing disorders. Recent advancements in artificial intelligence have driven the development of automated sleep staging models, which still face two significant challenges. 1) Simultaneously extracting prominent temporal and spatial sleep features from multi-channel raw signals, including characteristic sleep waveforms and salient spatial brain networks. 2) Capturing the spatial-temporal coupling patterns essential for accurate sleep staging. To address these challenges, we propose a novel framework named ST-USleepNet, comprising a spatial-temporal graph construction module (ST) and a U-shaped sleep network (USleepNet). The ST module converts raw signals into a spatial-temporal graph to model spatial-temporal couplings. The USleepNet utilizes a U-shaped structure originally designed for image segmentation. Similar to how image segmentation isolates significant targets, when applied to both raw sleep signals and ST module-generated graph data, USleepNet segments these inputs to extract prominent temporal and spatial sleep features simultaneously. Testing on three datasets demonstrates that ST-USleepNet outperforms existing baselines, and model visualizations confirm its efficacy in extracting prominent sleep features and temporal-spatial coupling patterns across various sleep stages. The code is available at: https://github.com/Majy-Yuji/ST-USleepNet.git.
@@ 1 Introduction
Sleep staging is essential for assessing sleep quality and diagnosing sleep disorders. Traditionally, sleep staging relies on the analysis of polysomnography (PSG) data, which includes mainly electroencephalogram (EEG) and other physiological signals such as electrooculogram (EOG), electromyogram (EMG), and electrocardiogram (ECG). These signals are divided into 30-second epochs, which are manually classified by experts into five stages—Wake, N1, N2, N3, and REM—according to the American Academy of Sleep Medicine (AASM) standards [1]. However, manual sleep staging is a labor-intensive process, prone to variability and human error, and lacks scalability.
To enhance the efficiency and accuracy of sleep staging, researchers have turned to artificial intelligence to develop automated sleep staging models [2, 3, 4]. Despite considerable advancements, there are also two unresolved challenges in this field.
@@
Figure 1: (a) and (b) display characteristic sleep waveforms from different stages: (a) features spindle waves from N2, and (b) delta waves from N3. (c) presents brain networks and salient subnetworks during N2 and N3, with salient subnetworks formed by retaining only the green nodes. (d) illustrates different types of spatial-temporal relationships. Taking the green node as an example, its temporal correlations at different time points are shown by red arrows, its spatial dependence with other nodes at the same time by black arrows, and its spatial-temporal coupling with other nodes at different times by blue arrows.
@@
How to simultaneously capture the characteristic sleep waveforms and salient spatial brain networks from multi-channel raw signals? Different sleep stages exhibit prominent temporal and spatial features, mainly including the characteristic sleep waveforms and salient spatial brain networks as outlined by AASM sleep standards [1]. For instance, Figure 1 illustrates the prominent sleep features of the N2 and N3 sleep stages. Figures (a) and (b) display the characteristic sleep waveforms over time, with the N2 stage marked by spindles and the N3 stage by delta waves. Figure 1 (c) shows the brain networks during the N2 and N3 stages, highlighting the subnetworks—salient spatial brain networks extracted by retaining only some nodes (the green nodes in the figure). It is evident that the subnetworks during the N2 and N3 stages are significantly different; during the lighter sleep of N2, brain activity is more pronounced, resulting in denser and more extensive network connections [5]. Given the importance of extracting prominent sleep features, SalientSleepNet [6] effectively extracts salient waveforms from single-channel signals and has achieved notable results. However, the work on single-channel signals cannot be directly extended to multi-channel signals. That is because multi-channel signals not only exhibit characteristic sleep waveforms across each single channel, as shown in Figures 1 (a) and (b) with characteristic sleep waveforms simultaneously in two channels, but also, as the signals interact, they form different functional connections at various sleep stages [7], thereby creating complex spatial patterns. Consequently, simultaneously capturing both multi-channel characteristic sleep waveforms and salient spatial brain networks is crucial for sleep staging and presents a significant challenge.
@@
How to capture the spatial-temporal coupling patterns in sleep staging? Multiple brain regions exhibit complex yet crucial spatial-temporal coupling patterns as they evolve over time [8]. Figure 1 (c) illustrates the spatial-temporal relationships within the brain network over time. In the figure, each circle represents a PSG data channel, corresponding to electrode placement on the brain; each arrow signifies a connection within the brain network over time, with different arrows indicating different types of relations. Specifically, red arrows indicate temporal correlations within the same channel at different time points, while black arrows represent spatial dependence at the same time point across different channels. Given the complexity of brain network connections, there is also a spatial-temporal coupling feature between different channels at different times [9], indicated by blue arrows in the figure. Most existing models can only capture temporal correlations or spatial dependences. For instance, some CNN/RNN-based models [10, 11, 12, 13] can capture temporal correlations, while some GNN-based models [14, 15] can capture spatial dependences. Therefore, capturing these spatial-temporal coupling features for sleep staging presents a challenge.
@@
To address the above challenges, we propose ST-USleepNet, a raw-data-based novel spatial-temporal coupling prominence framework for sleep staging. It consists of two main components: a spatial-temporal graph construction module (ST) and a U-shaped sleep network (USleepNet). The ST module converts multi-channel raw signals into a spatial-temporal graph, allowing for the modeling and capturing of spatial-temporal couplings. On the other hand, the USleepNet, which leverages a U-shaped structure originally developed for image segmentation [16], segments the prominent temporal and spatial sleep features of signals data and the graph constructed by the ST like how image segmentation isolates the main target of an image. Therefore, the USleepNet has two streams: for the temporal stream, we utilize a fully convolutional U-Net to segment characteristic sleep waveforms; for the spatial stream, we employ a graph convolutional U-Net to isolate salient spatial brain networks. The key contributions of this work include the following:
@@
  • Experimental results indicate that our ST-USleepNet model reaches state-of-the-art performance across three sleep staging datasets. Model visualizations reveal that our method can effectively extract prominent sleep features within different sleep stages.
@@ 2 Related Work
Initially, traditional machine learning methods such as Random Forest (RF) [17] and Support Vector Machine (SVM) [18] [19] were employed for sleep staging. However, most of these models heavily rely on expert knowledge, requiring the extraction of hand-crafted features before undertaking the sleep staging task, making them more time-consuming and expensive. To overcome this, researchers have started leveraging deep learning to bypass the need for manual feature engineering, primarily convolutional neural networks (CNNs) [11] [20] [12] [13] and Recurrent Neural Networks (RNNs) [21], to automatically extract high-dimensional deep features. However, such methods cannot be trained end-to-end.
Therefore, recent efforts have increasingly focused on refining end-to-end architectures that can efficiently process temporal correlations in sleep data without the preliminary steps required. Researchers mainly use CNNs [3] [6], RNNs [10], or a combination of both [2] [4], or Transformer [22] to model the temporal information for sleep staging. Perslev et al. [3] introduced the U-time model, which uses a CNN-based Unet to model single-channel signals. Jia et al. [6] proposed SalientSleepNet, which uses a $U^{2}$-structured CNN Unet combined with an attention mechanism to identify typical waveforms in single sleep signals. Phan et al. [22] used a transformer model and utilized attention scores to visualize sleep-related temporal features and the influence of different epochs. However, these methods mainly focus on capturing abstract temporal features instead of the Characteristic sleep waveforms across multiple channels, leading to limited accuracy.
However, the performance of temporal models suffers when modeling multi-channel signals because they ignore the spatial relations among brain regions, which are critically important [23]. Therefore, researchers have used Graph Neural Networks (GNNs) and Transformer models in sleep staging to capture spatial features. Jia et al. proposed GraphSleepNet [14], the first to apply a spatial-temporal graph convolution network (ST-GCN) to the sleep staging task. Subsequently, researchers introduced more graph-based methods. For example, Jia et al. proposed MSTGCN [20] and Ji et al. proposed JK-STGCN [15]. Both methods successfully modeled the spatial characteristics of the brain. Liu et al. proposed the BSTT [24] model, which uses a spatial-temporal transformer to capture spatial brain features. However, these methods mainly focus on capturing global spatial features instead of the salient spatial brain networks and ignore the spatial-temporal coupling patterns.
To address the limitation in existing methods and simultaneously capture the characteristic sleep waveforms and salient spatial brain networks from multi-channel raw signals, and capture the spatial-temporal coupling patterns within different sleep stages, we introduce ST-USleepNet for sleep staging.
3 Preliminary
In this study, the input is a segment of multi-channel physiological signals $s^{C\times f_{s}T}$, where $C$ represents the number of channels, $f_{s}$ denotes the sampling frequency, and $T$ indicates the duration of each segment. In sleep staging, it is common to have $T=30\,\text{s}$. The sleep staging problem is defined as the task of training a neural network $F$, to map a segment of multi-channel physiological signal data $s^{C\times f_{s}T}$, to a specific sleep stage $\hat{y}$. According to the AASM standard [1], each $\hat{y}\in{0,1,2,3,4}$ is matched to five sleep stages Wake, N1, N2, N3, and REM.
@@ 4 Methodology
Figure 2: The figure illustrates the overall framework in the central gray box while surrounding elements detail network layers and provide a simple example. There are two channels of one segment physiological signal in the upper left, represented in green and blue, as an example input. If the segment is viewed as three time intervals, then the input signals will be cut into 6 slices. The spatial-temporal graph construction module then forms a six-node graph and inputs it into encoders, where temporal and graph encoders downsample the feature of nodes (the initial feature is the raw slice signal) and graph respectively, yielding a subgraph with essential features. Taking this subgraph as input, the decoder includes a spatial decoder that restores the graph to a supergraph, and a temporal encoder that restores the graph’s feature dimensions, ultimately producing a feature graph identical in dimension to the original spatial-temporal graph. By concatenating the feature matrix of this graph according to the time and channel information, point-wise and segment-wise labels are generated.
The framework of ST-USleepNet is illustrated in Figure 2 and it consists of two primary modules. Initially, a graph construction module is developed to capture the spatial-temporal coupling patterns, converting raw signals into a spatial-temporal graph. Subsequently, we introduce USleepNet, which is designed to extract characteristic sleep waveforms and identify spatial salient brain networks from these signals. USleepNet is composed of two intertwined streams: the temporal stream, also called the temporal prominence network, represented by yellow blocks in Figure 2, and the spatial stream, also called the spatial prominence network, depicted by orange blocks. Each stream comprises $l$ encoder and $l$ decoder blocks, with $l=2$ exemplified in the figure. Additionally, the temporal stream integrates a segment classifier. To effectively combine the two streams, a fusion strategy is implemented.
@@ 4.1 Spatial-Temporal Graph Construction
The spatial-temporal graph construction module aims to transform the raw signals into a spatial-temporal graph. Given a multi-channel physiological signal data $s^{C\times f_{s}T}$, to fully capture the spatial-temporal coupling patterns within the segments, we cut the data $s^{C\times f_{s}T}$ into $n$ slices, resulting in $X^{nC\times f_{s}t}$. We use the raw signals $X$ directly as the initial features of the graph, then we have $G_{s}=(V,A^{nC\times nC},X^{nC\times f_{s}t})$, where $V$ represents the set of nodes, $A$ is the adjacency matrix, and $X$ is the features of nodes. The nodes of the graph $G_{s}$ are constituted by various slice data: node $N_{c_{i}t_{j}}$ represents a channel $c_{i}$, where $i\in\{k\in\mathbb{Z}\mid 0\leq k<C\}$, at a certain time point $t_{j}$, where $j\in\{k\in\mathbb{Z}\mid 0\leq k<n\}$. The feature corresponding to node $V_{c_{i}t_{j}}$ in $G_{s}$ is the raw signal of the respective slice ${X_{ij}}^{1\times f_{s}t}$. Next, we will discuss the construction of the adjacency matrix $A^{nC\times nC}$.
@@
The adjacency matrix $A$ is derived from the multiplication of a similarity matrix with several weighted masks. First, an adjacency matrix is established based on the similarity between different features. This is because highly similar signals often indicate that they share similar or identical functional states. Cosine similarity is commonly used to quantitatively measure the similarity between two physiological signals:
$sim\_cos(X_{i},X_{j})=\frac{X_{i}\cdot X_{j}}{\sqrt{\sum_{k=1}^{n}X_{ik}^{2}}\sqrt{\sum_{k=1}^{n}X_{jk}^{2}}}$ (1)
where $X_{i}$ and $X_{j}$ represent features of two different nodes, respectively. However, not all nodes need to be assigned a weight, and establishing weighted edges between features with minimal similarity could negatively impact model training. Thus, a threshold $\delta$ is set, where an edge is only established between two nodes if their feature similarity exceeds $\delta$. Consequently, the similarity matrix $S$ can be established as follows:
$s_{ij}=\begin{cases}sim\_cos(X_{i},X_{j})&\text{if }sim\_cos(X_{i},X_{j})>\delta\\ 0&\text{otherwise}\end{cases}$ (2)
@@
Next, we add some prior knowledge. One is to add the temporal information: slices closer in time should have a higher similarity. Therefore, a temporal weighted mask $M_{t}$ based on the time interval between two slices is constructed. Given a hyperparameter $\delta_{t}\in(0,1)$, the temporal weight ${m_{t}}_{ij}$ between the feature $X_{c_{i}}{t_{i}}$ and $X_{c_{j}}{t_{j}}$ can be expressed as:
${m_{t}}_{ij}=\delta_{t}^{|t_{i}-t_{j}|}$ (3)
@@
Similarly, it is also necessary to consider the positional features of channels between slices. The similarity between slice signals from different channels should be lower than that between slices from the same channel. Therefore, define hyperparameters $\delta_{p}\in(0,1)$ to penalize the edge weights of features from different channels. Thus, the positional weight ${m_{p}}_{ij}$ between the feature $X_{c_{i}}{t_{i}}$ and $X_{c_{j}}{t_{j}}$ can be expressed as:
${m_{p}}_{ij}=\begin{cases}\delta_{p}&\text{if }c_{i}\neq c_{j}\\ 1&\text{otherwise}\end{cases}$ (4)
Therefore, the adjacency matrix $A$ for $G_{s}$ can be represented as:
$A=S\cdot M_{t}\cdot M_{p}$ (5)
@@
According to the above rules, a segment of the original multi-channel physiological signal data $s^{C\times f_{s}T}$ can be transformed into a spatial-temporal graph $G_{s}=(V,A^{nC\times nC},X^{nC\times f_{s}t})$.

### v2

@@ Abstract
Sleep staging is critical to assess sleep quality and diagnose disorders. Despite advancements in artificial intelligence enabling automated sleep staging, significant challenges remain: (1) Simultaneously extracting prominent temporal and spatial sleep features from multi-channel raw signals, including characteristic sleep waveforms and salient spatial brain networks. (2) Capturing the spatial-temporal coupling patterns essential for accurate sleep staging. To address these challenges, we propose a novel framework named ST-USleepNet, comprising a spatial-temporal graph construction module (ST) and a U-shaped sleep network (USleepNet). The ST module converts raw signals into a spatial-temporal graph based on signal similarity, temporal, and spatial relationships to model spatial-temporal coupling patterns. The USleepNet employs a U-shaped structure for both the temporal and spatial streams, mirroring its original use in image segmentation to isolate significant targets. Applied to raw sleep signals and graph data from the ST module, USleepNet effectively segments these inputs, simultaneously extracting prominent temporal and spatial sleep features. Testing on three datasets demonstrates that ST-USleepNet outperforms existing baselines, and model visualizations confirm its efficacy in extracting prominent sleep features and temporal-spatial coupling patterns across various sleep stages. The code is available at: https://github.com/Majy-Yuji/ST-USleepNet.git.
@@ 1 Introduction
Sleep disorders are associated with a wide range of health issues [1], including cardiovascular diseases [2], metabolic disorders [3], and neurodegenerative conditions [4]. Accurate diagnosis and management of these disorders heavily rely on sleep staging [5], a process that involves categorizing sleep into distinct stages based on physiological signals. This classification provides essential insights into sleep patterns and plays a critical role in identifying abnormalities.
@@
Figure 1: An illustration of brain activities during different sleep stages. (a) and (b) display characteristic sleep waveforms from different stages. (c) presents brain networks and salient subnetworks during N2 and N3. (d) illustrates different types of spatial-temporal relationships.
Sleep staging is traditionally performed using polysomnography (PSG), which primarily analyzes electroencephalogram (EEG) signals, but also incorporates other physiological measurements such as electrooculogram (EOG) and electromyogram (EMG). These signals are segmented into 30-second epochs and manually classified by experts into standard stages, including wake, non-REM (N1, N2, N3) sleep stages and rapid eye movement (REM) [6]. However, this manual process is labor-intensive and challenging to scale for large datasets [7]. To improve the efficiency and accuracy of sleep staging, researchers have turned to artificial intelligence to develop automated sleep staging models [8, 9, 10]. Despite advances, there are two unresolved challenges in this field.
@@
How to simultaneously capture the characteristic sleep waveforms and salient spatial brain networks from multi-channel raw signals? Different sleep stages exhibit prominent temporal and spatial features, primarily the characteristic sleep waveforms and salient spatial brain networks as outlined by the AASM sleep standards [6]. Figures 1 (a) and (b) show characteristic sleep waveforms over time. Panel (a) features spindle waves from the N2 stage, while panel (b) displays delta waves from the N3 stage. Figure 1 (c) presents the brain networks and salient spatial subnetworks during both N2 and N3 stages, with the salient subnetworks highlighted by retaining only the green nodes. Given the crucial information contained in these prominent sleep features, many studies [11, 12, 13] have effectively extracted salient waveforms from sleep signals. However, these approaches are limited to single-channel signals, whereas PSG signals are predominantly multi-channel. The prominent sleep features in multi-channel signals are significantly more complex than those in single-channel signals. This complexity has two main sources. One arises from the characteristic sleep waveforms observed across different channels, as illustrated in Figures 1 (a) and (b), which show simultaneous waveforms in two channels. The other source comes from the formation of distinct functional connections at various sleep stages due to interactions between signals [14, 15]. These interactions result in intricate spatial patterns that further enrich the complexity of multi-channel signals. Consequently, simultaneously capturing both multi-channel characteristic sleep waveforms and salient spatial brain networks is crucial for accurate sleep staging and presents a significant challenge.
@@
How to capture the spatial-temporal coupling patterns to better understand brain activities? Signals from different brain regions exhibit complex yet crucial spatial-temporal coupling patterns as they evolve over time [16]. Figure 1 (d) illustrates the spatial-temporal relationships within the brain network over time. In the figure, each circle represents a signal channel, corresponding to an electrode placement on the brain; each arrow signifies a connection within the brain network over time. Using the green node as an example, different arrows indicate distinct types of relationships: red arrows denote temporal correlations within the same channel across different time steps, black arrows symbolize spatial dependencies at the same time step across various channels, and blue arrows represent spatial-temporal coupling patterns between different channels at different times. These spatial-temporal connections within the brain are crucial for revealing brain activities [17]. Most existing models can only capture temporal correlations [18, 19, 20, 21] or spatial dependencies [22, 23]. Therefore, capturing these spatial-temporal coupling patterns for sleep staging presents a challenge.
@@
To address the above challenges, we propose ST-USleepNet, a raw-signal-based novel spatial-temporal coupling prominence framework for sleep staging. It consists of two main components: a spatial-temporal graph construction module (ST) and a U-shaped sleep network (USleepNet). The ST module converts multi-channel raw signals into a spatial-temporal graph, allowing for the modeling and capturing of spatial-temporal coupling patterns. On the other hand, the USleepNet, which leverages a U-shaped structure originally developed for image segmentation [24], segments the signals and the graph constructed by the ST to obtain prominent temporal and spatial sleep features like how image segmentation isolates the main target of an image. Therefore, the USleepNet has two streams: for the temporal stream, we utilize a fully convolutional U-Net to segment characteristic sleep waveforms; for the spatial stream, we employ a graph convolutional U-Net to isolate salient spatial brain networks. The key contributions of this work include the following:
@@
  • Experimental results indicate that our ST-USleepNet model reaches state-of-the-art performance across three sleep staging datasets.
  • Model visualizations demonstrate that our method effectively extracts prominent sleep features and spatial-temporal coupling patterns across different sleep stages, providing a certain degree of interpretability for the method.
@@ 2 Related Work
Traditional sleep staging relied on machine learning methods like Random Forest (RF) [25] and Support Vector Machine (SVM) [26, 27], which required manually crafted features based on expert knowledge, making the process costly and inefficient. To overcome this, deep learning approaches such as convolutional neural networks (CNNs) [19, 28, 20, 21] and Recurrent Neural Networks (RNNs) [29] were introduced to automatically extract high-dimensional features. However, these methods still follow a two-stage pipeline and lack end-to-end training capabilities.
Recent efforts have focused on end-to-end architectures for processing temporal correlations in sleep staging without preliminary steps. Temporal modeling approaches include CNNs [9, 11, 30], RNNs [18], their combinations [8, 10], and Transformers [31]. For instance, U-time [9] employed a CNN-based Unet for single-channel signals, while SalientSleepNet [11] integrated a $U^{2}$-structured CNN Unet with attention to identify typical waveforms. Some studies have also focused on extracting typical waveforms to perform expert-like classification [12, 13]. Despite their advancements, these methods primarily focus on single-channel or abstract temporal features, overlooking spatial dependencies.
Spatial dependencies among brain regions are critically important [32]. GraphSleepNet [22] was the first to apply a spatial-temporal graph convolution network (ST-GCN) for sleep staging. Subsequently, researchers introduced more graph-based methods [28, 23]. Besides, Liu et al. proposed the BSTT [33] model, which uses a spatial-temporal transformer to capture spatial brain features. However, these spatial methods mainly focus on capturing global spatial features instead of the salient spatial brain networks and ignore the spatial-temporal coupling patterns.
To address the limitations of existing methods and effectively capture characteristic sleep waveforms, salient spatial brain networks, and spatial-temporal coupling patterns across different sleep stages from multi-channel raw signals, we introduce ST-USleepNet for sleep staging.
@@ 3 Methodology
Figure 2: An illustration of ST-USleepNet. It shows the overall framework within the central gray box, while the surrounding elements detail the network layers and provide a simple example using signals from two channels.
The framework of ST-USleepNet is illustrated in Figure 2 and it consists of two primary modules. First, a spatial-temporal graph construction module is developed to capture the spatial-temporal coupling patterns, converting raw signals into a spatial-temporal graph. Next, we introduce USleepNet, which is designed to extract characteristic sleep waveforms and identify spatial salient brain networks from these signals. USleepNet is composed of two interwoven streams: the temporal stream, also called the temporal prominence network, represented by yellow blocks in Figure 2, and the spatial stream, also called the spatial prominence network, depicted by orange blocks. Each stream comprises $l$ encoder and $l$ decoder blocks, with $l=2$ exemplified in the figure. Additionally, the temporal stream integrates a segment classifier. To effectively combine the two streams, a fusion strategy is implemented.
@@ 3.1 Spatial-Temporal Graph Construction
The spatial-temporal graph construction module aims to transform the raw signals into a spatial-temporal graph. Given a multi-channel physiological signal data $s^{C\times f_{s}T}$, where $C$ represents the number of channels, $f_{s}$ denotes the sampling frequency, and $T$ indicates the duration of each segment, the data $s^{C\times f_{s}T}$ is divided into $n$ patches, resulting in $X^{nC\times f_{s}t}$. We use the raw signals $X$ directly as the initial features of the graph, then we have $G_{s}=(V,A^{nC\times nC},X^{nC\times f_{s}t})$, where $V$ represents the set of nodes, $A$ is the adjacency matrix. The nodes of the graph $G_{s}$ are constituted by patches: node $N_{c_{i}t_{j}}$ represents a channel $c_{i}$, where $i\in\{k\in\mathbb{Z}\mid 0\leq k<C\}$, at a certain time step $t_{j}$, where $j\in\{k\in\mathbb{Z}\mid 0\leq k<n\}$. Next, we will discuss how to construct the adjacency matrix $A^{nC\times nC}$.
Figure 3: An example of spatial-temporal graph construction with a 2-channel signal divided into 3 patches per channel. Each patch represents a node, and the adjacency matrix is obtained by multiplying the similarity matrix, temporal weighted mask, and positional weighted mask.
@@
The adjacency matrix $A$ is derived from the multiplication of a similarity matrix with several weighted masks, as shown in Figure 3. First, an adjacency matrix is established based on the cosine similarity between different features, as highly similar signals often indicate similar functional states. To avoid assigning weights to unrelated nodes, a threshold $\delta$ is applied, establishing edges only between nodes with similarity exceeding $\delta$. The similarity matrix $S$ is then defined as follows:
$s_{ij}=\begin{cases}sim\_cos(X_{i},X_{j})&\text{if }sim\_cos(X_{i},X_{j})>\delta\text{,}\\ 0&\text{otherwise}\text{.}\end{cases}$ (1)
@@
Here, $X_{i}$ and $X_{j}$ represent features of two different nodes; $sim\_cos$ represents the cosine similarity function. Next, we add some prior knowledge using other weighted masks. One is to add the temporal information: patches closer in time should have higher similarity. A temporal weighted mask $Mt$ is constructed based on the time interval between patches. With a hyperparameter $\delta_{t}\in(0,1)$, the temporal weight ${mt}_{ij}$ between the feature $X_{c_{i}t_{i}}$ and $X_{c_{j}t_{j}}$ is defined as:
${mt}_{ij}=\delta_{t}^{|t_{i}-t_{j}|}\text{,}$ (2)
where $t_{i}$ and $t_{j}$ are the time steps of the patches, and $t_{i},t_{j}\in\{k\in\mathbb{Z}\mid 0\leq k<n\}$.
@@
Similarly, positional features between channels should be considered because patches from different channels should have lower similarity than those from the same channel. A hyperparameter $\delta_{p}\in(0,1)$ is defined to penalize edge weights between different channels, forming a positional weighted mask $Mp$. Thus, the positional weight ${mp}_{ij}$ between the feature $X_{c_{i}t_{i}}$ and $X_{c_{j}t_{j}}$ can be expressed as:
${mp}_{ij}=\begin{cases}\delta_{p}&\text{if }c_{i}\neq c_{j}\text{,}\\ 1&\text{otherwise}\text{.}\end{cases}$ (3)
Therefore, the adjacency matrix $A$ for $G_{s}$ is:
$A=S\cdot Mt\cdot Mp\text{.}$ (4)
@@
According to the above rules, the multi-channel signal segment $s^{C\times f_{s}T}$ is transformed into a spatial-temporal graph $G_{s}=(V,A^{nC\times nC},X^{nC\times f_{s}t})$.